# Concept-Guided Interpretability via Neural Chunking

### v1

@@ Abstract
Neural networks are often described as black boxes, reflecting the significant challenge of understanding their internal workings and interactions. We propose a different perspective that challenges the prevailing view: rather than being inscrutable, neural networks exhibit patterns in their raw population activity that mirror regularities in the training data. We refer to this as the Reflection Hypothesis and provide evidence for this phenomenon in both simple recurrent neural networks (RNNs) and complex large language models (LLMs). Building on this insight, we propose to leverage cognitively-inspired methods of chunking to segment high-dimensional neural population dynamics into interpretable units that reflect underlying concepts. We propose three methods to extract these emerging entities, complementing each other based on label availability and neural data dimensionality. Discrete sequence chunking (DSC) creates a dictionary of entities in a lower-dimensional neural space; population averaging (PA) extracts recurring entities that correspond to known labels; and unsupervised chunk discovery (UCD) can be used when labels are absent. We demonstrate the effectiveness of these methods in extracting entities across varying model sizes, ranging from inducing compositionality in RNNs to uncovering recurring neural population states in large language models with diverse architectures, and illustrate their advantage to other interpretability methods. Throughout, we observe a robust correspondence between the extracted entities and concrete or abstract concepts in the sequence. Artificially inducing the extracted entities in neural populations effectively alters the network’s generation of associated concepts. Our work points to a new direction for interpretability, one that harnesses both cognitive principles and the structure of naturalistic data to reveal the hidden computations of complex learning systems, gradually transforming them from black boxes into systems we can begin to understand.
@@ 1 Introduction
Neural networks are sometimes described as “black box systems” templeton-2021-word ; petsiuk_rise_2018 ; ribeiro_why_2016-1 ; zhang_visual_2018 ; marks_sparse_2024 ; elhage2022toy - highlighting the inherent challenge to make sense of the vast number of computing units. Their computations are performed by up to billions of interacting components, far exceeding the conventional approach to manipulating well-defined symbolic entities in physics and math to understand the natural world.
@@
A substantial portion of the interpretability challenge is cognitive adadiberrada ; miller2018explanationartificialintelligenceinsights and demands asking: what makes high-dimensional data meaningful for people? And what constitutes the basic units and entities for perception, that underlie the constitution of the symbols that we are adept at studying?
@@
We present a novel perspective on interpreting artificial neural networks by referencing how entities arise from sensory data. Similar to the dimensionality of neural activities, perceptual data that flood into our sensory stream is also high-dimensional. We make sense of this “booming, buzzing confusion” by discerning recurring patterns in perceptual sequences as chunks and entities. Across almost every sensory modality, cognition parses high-dimensional sensory information as wholes. We inherently and rapidly identify recurring patterns in sequences with minimal exposure saffran_statistical_1996 ; saffran1997incidental ; aslin1998computation ; and decompose high-dimensional visual perception into recurring ’chunks’ of identities to perceive and recall (zacks2001event ; zacks2007event ; avrahami1994emergence ; brent1999efficient . In this way, the overwhelming flow of the sensory stream is converted into a structured representation of entities, relations, and events over timeGraybiel1998TheRepertoires ; Gobet2001ChunkingLearning ; Egan1979ChunkingDrawings ; Ellis1996SequencingOrder ; koch_patterns_2000 ; Chase&Simon1973 ; wu_chunking_2023 ; wu2025motifs .
@@
Psychologists have long postulated that identifying patterns in perception serves as an inverse model to uncover structured reality, in which recurring components often reflect shared causal mechanisms newell1972human ; barker1963stream . We termed this the "reflection hypothesis" and examined if the neural network’s hidden activity reflects the regularities in the data it learns to predict, first in RNNs trained to predict sequences with known regularities, and then in large language models. Evidence that structured information exists in high-dimensional neural data suggests the possibility of leveraging the cognitive tendency to chunk information and discern recurring patterns and entities directly from neural population activities. Hence, we developed several methods to extract these recurring entities from low to high dimensions of neural population activities, complementing each other in their usage cases. We show the advantage of this simple method over popular interpretability approaches. Our findings suggest a novel interpretability framework that leverages the cognitive principle of chunking to find emerging entities within artificial neural activities, paving the way for future work that decomposes network computation into interactions and relations between symbolic entities.
@@ 2 Related Work
Current methods on what to interpret to understand the computations inside a neural network are heavily influenced by neuroscience: either on the level of neurons as a computation unit or in low-dimensional neural activity descriptions. The earliest interpretability approaches, inspired by discoveries of “grandmother cells” and “Jennifer Aniston neurons” inside the brain, focused on understanding the semantic meanings that drive the activity of individual neurons. Similarly, studies in ANNs, from BERT to GPT, have identified specific neurons and attention heads whose activations correlate with semantic meanings in the data olah2020neuron ; elhage2022toy ; wang2022interpretabilitywildcircuitindirect ; marks2024sparsefeaturecircuitsdiscovering ; bau2020concept ; goh2021multimodal ; nguyen2016synthesizing ; mu2021compositionalexplanationsneurons ; radford2017learninggeneratereviewsdiscovering . Sparse autoencoders (SAEs)can be seen as an intermediate step that encourage the hidden neurons to be more monosemantic gao2024scaling ; makhzani2013k ; cunningham2023sparseautoencodershighlyinterpretable ; chaudhary2024evaluatingopensourcesparseautoencoders ; karvonen2024evaluatingsparseautoencoderstargeted . Thereby, an autoencoder maps neural activities of a hidden unit layer to a much larger number of intermediate hidden units while encouraging a small number of them to be active. This way, the activity of the target hidden layer can be represented as a superposition of several individual neurons within the SAE. Other approaches such as representation engineering captures the distinct neural activity corresponding to the target concept or function, such as bias or truthfulness zou_representation_2023 . Then, it uses a linear model to identify the neural activity direction that predicts the concept under question or interferes with the network behavior.
@@ 3 Chunks from Neural Embeddings (CNE)
Similar to humans, AI systems aim to understand reality from naturalistic data. The regularities in naturalistic data may drive converging representations in diverse AI models. Past research indicates that neural networks of different architectures, scales, and sizes learn representations that are remarkably similar balestriero18b ; moschella2023relativerepresentationsenablezeroshot , even when trained on different data sources lenc2015understandingimagerepresentationsmeasuring . This phenomenon is particularly pronounced in larger, more robust models bansal21 ; dravid2023rosettaneuronsminingcommon ; kornblith19a ; roeder21a , across multiple data modalities lenc2015understandingimagerepresentationsmeasuring ; moschella2023relativerepresentationsenablezeroshot ; kornblith19a ; roeder21a . Recently, it has been hypothesized that such convergences reflect an underlying statistical model of reality, aligning with a Platonic conceptual framework huh2024platonicrepresentationhypothesis .
The Reflection Hypothesis: We propose that artificial neural networks, like human cognition, may reflect the regularities and redundancies present in the data they are trained on within their neural computations. Specifically, we hypothesize that a well-trained neural network should exhibit trajectories of neural computation that mirror the structure of its training data.
If this hypothesis holds, it could allow us to leverage the human cognitive ability to “chunk” high-dimensional perceptual data into meaningful units. Just as human perception breaks down complex sensory inputs into recurring, segregated patterns, neural network activities might also be decomposable into recurring "chunks". These chunks, residing within subsets of all neurons of the network, could represent distinct patterns of computation and serve as interpretable entities. By identifying and analyzing these chunks, we could reduce the complexity of neural activities into interactions between these interpretable units.
To test this hypothesis, we develop and explore methods to extract these chunks across a spectrum of settings, starting from simple models trained on straightforward sequences to complex models trained on language. This approach aims to bridge the gap between human cognition and the comprehension of high-dimensional neural activities.
@@ 3.1 Extracting Entities
Formally, denote the training sequence as $S=(s^{1},s^{2},\dots,s^{n})$, indexed by $I=\{1,2,3,\cdots,n\}$. A neural network with a set of neurons $W$ ($|W|=d$) exerts a sequence of population activities $S_{h}=(\mathbf{h}^{1},\mathbf{h}^{2},\cdots,\mathbf{h}^{n})$ in response to the input sequence. Each neural population vector has embedding dimension $d$: $\mathbf{h}^{i}\in\mathbb{R}^{d}$. Let $C\subseteq W$ be a subset of neurons. The neural embedding space is a subspace of $\mathbb{R}^{d}$ which includes possible neural activities elicited by the network. A chunk over $C$ is defined as a closed ball $\overline{B}(\overline{\mathbf{h}_{C}},\Delta)$, centered at the mean activity $\overline{\mathbf{h}_{C}}$, such that a neural activity vector $\mathbf{h}$ belongs to the chunk if $\frac{\|\mathbf{h}_{C}-\overline{\mathbf{h}_{C}}\|_{2}^{2}}{d}\leq\Delta$. A chunk is a frequently revisited subspace of the embedding space localized to a subset of neurons.
@@
When $d$ is relatively small, we can use a cognitive-inspired method to extract recurring patterns wu_learning_2022 , which is also used for text compression gage1994new ; zaki2000sequence ; agrawal1995mining or tokenization. The idea is to cluster the neuron activations individually, use the cluster indices to represent the network activity as strings, and then further group frequently occurring string sequences into chunks. For this purpose, we convert each hidden neural activity vector $\mathbf{h}\in\mathbb{R}^{d}$ into a string of discrete integers. In this way, the sequence of vectors $S_{h}$ is transformed into a one-dimensional sequence of strings $S^{s}=(l^{1},l^{2},...,l^{n})$. Each $l^{j}$ has length $d$ and contains the nearest cluster index for each of the $d$ neurons, i.e.,
$l^{j}=\text{concat}(\mathcal{M}(h_{1}),\mathcal{M}(h_{2}),\dots,\mathcal{M}(h_{d}))$
$\mathcal{M}$ is a clustering function that assigns the closest cluster index to the activity of the neuron.
We then apply a chunking method on $S^{s}$ to extract a vocabulary of patterns of string combinations, which represent patterns in neural state trajectories. The vocabulary of chunks $\mathbf{D}$ is initialized with the set of unique strings in $S_{s}$.This vocabulary grows iteratively by merging the top $N$ frequent adjacent chunk pairs above a threshold. Within each new iteration, the updated vocabulary is used to reparse the sequence, repeating until convergence (see SI 1). This identifies recurring neural population trajectories that segment the activity sequence. The resulting dictionary $\mathbf{D}$ can then parse $S^{s}$ in terms of extracted chunks.
@@
In larger-scale networks like transformers, $d$ can be very big making clustering individual activations with DSC computationally expensive. Moreover, complex sequences, e.g. natural language, exhibit patterns at both surface and more abstract levels, such as words or grammatical categories. In this scenario, identifying recurring patterns in the sequence allows us to uncover the internal structures of neural population chunks by averaging the relevant population activities, akin to extracting task-induced neural response functions Cohen2011 ; Engel2001 ; Churchland2012 .
@@
, where $\boldsymbol{\mu}=\mathbb{E}[\mathbf{h}_{C(s)}]$ is the true mean of the subpopulation neural activity.
To do this, we first compute the mean population response of signal $s$ by averaging over the token-specific hidden state representations:
$\overline{\mathbf{h}}=\frac{\sum_{j\in V(s)}\mathbf{h}_{(j)}}{|V(s)|}$
A neuron $i$ is hypothesized to be inside ${C(s)}$ if its activity at the time of the concept-specific index fluctuates within a pre-set tolerance level around the mean signal-relevant activity
$C(s)=\{i\in W:|h_{i,j}-\overline{h_{i}}|\leq\text{tol}\quad\forall j\in V(s)\}$
$\Delta=\max_{j\in V(s)}\frac{||\mathbf{h}_{{C(s),j}}-\overline{\mathbf{h}_{C(s)}}||_{2}^{2}}{d}$ (2)
@@
and evaluate its quality as a neural population activity classifier for the occurrence of the signal $s$. This evaluation is based on the true positive rate (TPR) and false positive rate (FPR).
@@
This method relies on prior knowledge of recurring patterns $s$ in the input sequence and assumes that the network’s activity remains within a small deviation threshold around the mean chunk activity vector, $\overline{\mathbf{h}_{C(s)}}$. If the network generates multiple chunks to distinguish contextual variations, the identifiability of the concept-representing recurring activity inside the network may decrease due to the averaging, potentially limiting its utility in highly dynamic contexts.
@@ 4.1 Evaluating Reflection Hypothesis on Simple RNNs
First, we assessed the reflection hypothesis using a simple recurrent neural network (Figure 2a and section A.1) trained on synthetic sequences containing a known pattern. We begin with sequences in which ABCD appear periodically (Figure 2b), then gradually increase complexity: first by embedding ABCD sparsely within a background of repeated Es (Figure 2c), and then by placing ABCD amid randomly occurring background symbols (E, F, G) to simulate noise (Figure 2d). In each case, the RNN is trained to predict the next character. Across all conditions, the network’s hidden states consistently exhibit activity patterns that reflect the pattern occurrences in the sequence.
@@
We apply Discrete Sequence Chunking to transform continuous hidden unit activities to a symbolic and discretized description. We denote population activities by the alternation of indices that mark the belonging cluster centroid of individual neurons and visualize the population activity by the indices of the assigned nearest cluster. This symbolic description of the neural activities allows decoding the trajectory of the hidden state and its corresponding input via a look-up table (section A.2), and reaches a perfect decoding accuracy on the test set. From the symbolic description of neural trajectories, we can then apply chunk discovery methods to learn a dictionary of symbolized chunks. This dictionary contains the maximally recurring patterns inside the network, which consistently reflect the patterns in the input.
@@
Starting from a state predicting B from input A (with prior input E), we replace the hidden state with the one corresponding to A, B, or C, while pairing with inputs B, C, or D, reliably shifts the output predictably to C, D, or E (Figure 2 a).

### v2

@@ Abstract
Neural networks are often described as black boxes, reflecting the significant challenge of understanding their internal workings and interactions. We propose a different perspective that challenges the prevailing view: rather than being inscrutable, neural networks exhibit patterns in their raw population activity that mirror regularities in the training data. We refer to this as the Reflection Hypothesis and provide evidence for this phenomenon in both simple recurrent neural networks (RNNs) and complex large language models (LLMs). Building on this insight, we propose to leverage our cognitive tendency of chunking to segment high-dimensional neural population dynamics into interpretable units that reflect underlying concepts. We propose three methods to extract recurring chunks on a neural population level, complementing each other based on label availability and neural data dimensionality. Discrete sequence chunking (DSC) learns a dictionary of entities in a lower-dimensional neural space; population averaging (PA) extracts recurring entities that correspond to known labels; and unsupervised chunk discovery (UCD) can be used when labels are absent. We demonstrate the effectiveness of these methods in extracting concept-encoding entities agnostic to model architectures. These concepts can be both concrete (words), abstract (POS tags), or structural (narrative schema). Additionally, we show that extracted chunks play a causal role in network behavior, as grafting them leads to controlled and predictable changes in the model’s behavior. Our work points to a new direction for interpretability, one that harnesses both cognitive principles and the structure of naturalistic data to reveal the hidden computations of complex learning systems, gradually transforming them from black boxes into systems we can begin to understand.
@@ 1 Introduction
Neural networks are known as “black boxes” templeton-2021-word ; petsiuk_rise_2018 ; ribeiro_why_2016-1 ; zhang_visual_2018 ; marks_sparse_2024 ; elhage2022toy , highlighting the inherent challenge to make sense of a vast number of computing units. Their computations are performed by up to billions of interacting components, in stark contrast to the traditional scientific practice of modeling phenomena through a number of well-defined symbolic entities in physics or mathematics.
@@
A substantial portion of the interpretability challenge is cognitive adadiberrada ; miller2018explanationartificialintelligenceinsights and demands asking the question: what makes high-dimensional data meaningful for cognition? And what constitutes the basic units and entities for perception that underlie the constitution of the symbols that we are adept at studying in physics or math?
@@
We present a novel perspective on interpreting artificial neural networks by referencing how cognitive entities arise from sensory data. Similar to the vast neural activities, perceptual data that flood into our sensory stream is similarly high-dimensional and complex. However, there is an inherent tendency for cognition to rapidly discern recurring patterns in perceptual sequences as chunks and entities (zacks2001event ; zacks2007event ; avrahami1994emergence ; brent1999efficient ) - a process that even infants do with minimal exposure saffran_statistical_1996 ; saffran1997incidental ; aslin1998computation . Our perceptual and cognitive system adeptly organizes the the overwhelming flow of the sensory stream into a structured representation of entities, relations, and events over time Graybiel1998TheRepertoires ; Gobet2001ChunkingLearning ; Egan1979ChunkingDrawings ; Ellis1996SequencingOrder ; koch_patterns_2000 ; Chase&Simon1973 ; wu_chunking_2023 ; wu2025motifs .
@@
Psychologists have long postulated that identifying patterns in perception serves as an inverse model to uncover structured reality, in which recurring elements often reflect shared causal mechanisms newell1972human ; barker1963stream . We termed this the "reflection hypothesis" and examined if the neural network’s hidden activity reflects the regularities in the data it learns to predict, first in RNNs trained to predict sequences with known regularities, and then in large language models. Evidence that structured information exists in high-dimensional neural data suggests the possibility of leveraging our cognitive tendency to chunk information and discern recurring patterns and entities directly from neural population activities. Hence, we developed several methods to extract these recurring entities from low to high dimensions of neural population activities, complementing each other in their usage cases. We show the advantage of this simple method over popular interpretability approaches. Our findings suggest a novel interpretability framework that leverages the cognitive principle of chunking to find emerging entities within artificial neural activities, paving the way for future work that decomposes network computation into interactions and relations between symbolic entities.
@@ 2 Related Work
Current methods on what to interpret to understand the computations inside a neural network are heavily influenced by neuroscience: either on the level of neurons as a computation unit or in low-dimensional neural activity descriptions. The earliest interpretability approaches, inspired by discoveries of “grandmother cells” and “Jennifer Aniston neurons” inside the brain, focused on understanding the semantic meanings that drive the activity of individual neurons. Similarly, studies in ANNs, from BERT to GPT, have identified specific neurons and attention heads whose activations correlate with semantic meanings in the data olah2020neuron ; elhage2022toy ; wang2022interpretabilitywildcircuitindirect ; marks2024sparsefeaturecircuitsdiscovering ; bau2020concept ; goh2021multimodal ; nguyen2016synthesizing ; mu2021compositionalexplanationsneurons ; radford2017learninggeneratereviewsdiscovering ; karpathy2015unreasonable . Sparse autoencoders (SAEs)can be seen as an intermediate step that encourage the hidden neurons to be more monosemantic gao2024scaling ; makhzani2013k ; cunningham2023sparseautoencodershighlyinterpretable ; chaudhary2024evaluatingopensourcesparseautoencoders ; karvonen2024evaluatingsparseautoencoderstargeted . Thereby, an autoencoder maps neural activities of a hidden unit layer to a much larger number of intermediate hidden units while encouraging a small number of them to be active. This way, the activity of the target hidden layer can be represented as a superposition of several individual neurons within the SAE. Other approaches such as representation engineering captures the distinct neural activity corresponding to the target concept or function, such as bias or truthfulness zou_representation_2023 . Then, it uses a linear model to identify the neural activity direction that predicts the concept under question or interferes with the network behavior.
@@ 3 Chunks from Neural Embeddings (CNE)
Similar to humans, AI systems aim to understand reality from naturalistic data. The regularities in naturalistic data may drive converging representations in diverse AI models. Recent findings suggests that neural networks of different architectures, scales, and sizes learn representations that are remarkably similar balestriero18b ; moschella2023relativerepresentationsenablezeroshot , even when trained on different data sources lenc2015understandingimagerepresentationsmeasuring . This phenomenon is more pronounced in larger, more robust models bansal21 ; dravid2023rosettaneuronsminingcommon ; kornblith19a ; roeder21a , and across multiple data modalities lenc2015understandingimagerepresentationsmeasuring ; moschella2023relativerepresentationsenablezeroshot ; kornblith19a ; roeder21a . Recently, it has been hypothesized that such convergences reflect an underlying statistical model of reality huh2024platonicrepresentationhypothesis ; engels2024notlinear .
The Reflection Hypothesis: We posit that the underlying cause of these observations is that, like human cognition, neural networks reflect the structures and regularities present in their training data through their internal computations. Specifically, we hypothesize that a well-trained neural network should exhibit trajectories of neural activity that mirror the structure of the data.
Formally, denote the sequence as $S=(s^{1},s^{2},\dots,s^{n})$, indexed by $I=\{1,2,3,\cdots,n\}$. A pattern in the input is defined as a recurring subsequence carrying the same concept $P=(s^{i},s^{i+1},\dots,s^{i+k-1})\subseteq S$ of fixed or variable length $k$. A neural network with a set of neurons $W$ ($|W|=d$) exerts a sequence of population activities $S_{h}=(\mathbf{h}^{1},\mathbf{h}^{2},\cdots,\mathbf{h}^{n})$ in response to the input sequence. Each neural population vector has embedding dimension $d$: $\mathbf{h}^{i}\in\mathbb{R}^{d}$.
The Reflection Hypothesis posits that such recurring input patterns $P$ are mirrored within the internal neural dynamics of the model, such that their corresponding neural activations $(\mathbf{h}^{i},\mathbf{h}^{i+1},\dots,\mathbf{h}^{i+k-1})$ lie within a population-level chunk defined as a ball $\overline{B}(\overline{\mathbf{h}}_{C},\Delta)\subset\mathbb{R}^{d}$ centered at a prototypical activation vector $\overline{\mathbf{h}}_{C}$ with radius $\Delta$. Regularities in the input sequence are reflected as localized activity patterns in subdimensions of neural population dynamics.
If this hypothesis is even partially true, it suggests that neural network activity can be decomposed into recurring “chunks,” analogous to how human perception segments high-dimensional visual input into meaningful units. Such chunks, elicited by activities of neural subpopulations, may represent distinct computations and serve as interpretable entities. Extracting these entities allows a reduction of the highly complex neural computation to the emergence of these chunks. We develop architecturally agnostic methods to extract chunks and evaluate their efficacy in concept encoding, interpretability, and their causal role in altering network behavior.
@@ 3.1 Extracting Entities
When $d$ is relatively small, we can use a cognitive-inspired method to extract recurring patterns wu_learning_2022 , which is also used for text compression gage1994new ; zaki2000sequence ; agrawal1995mining or tokenization. The idea is to cluster the neuron activations individually, use the cluster indices to represent the network activity as strings, and then further group frequently occurring string sequences into chunks. For this purpose, we convert each hidden neural activity vector $\mathbf{h}\in\mathbb{R}^{d}$ into a string of discrete integers, and transform the sequence of vectors $S_{h}$ into a one-dimensional sequence of strings $S^{s}=(l^{1},l^{2},...,l^{n})$. Each $l^{j}$ has length $d$ and contains the nearest cluster index for each of the $d$ neurons, i.e., $l^{j}=\text{concat}(\mathcal{M}(h_{1}),\mathcal{M}(h_{2}),\dots,\mathcal{M}(h_{d}))$. $\mathcal{M}$ is a clustering function that assigns the closest cluster index to the activity of the neuron.
We then apply a chunking method on $S^{s}$ to extract a vocabulary of patterns of string combinations, which represent patterns in neural state trajectories. The vocabulary of chunks $\mathbf{D}$ is initialized with the set of unique strings in $S^{s}$.This vocabulary grows iteratively by merging the top $N$ frequent adjacent chunk pairs above a threshold. Within each new iteration, the updated vocabulary is used to reparse the sequence, repeating until convergence (see SI 1). This identifies recurring neural population trajectories that segment the activity sequence. The resulting dictionary $\mathbf{D}$ can then parse $S^{s}$ in terms of extracted chunks.
@@
In larger-scale networks like transformers, $d$ can be very big.In this case, identifying recurring patterns in the sequence allows us to uncover the internal structures of neural population chunks by averaging the relevant population activities, akin to extracting task-induced neural response functions Cohen2011 ; Engel2001 ; Churchland2012 .
@@
$\overline{\mathbf{h}_{C(s)}}=\frac{\sum_{j\in V(s)}\mathbf{h}_{C(s),j}}{|V(s)|}\text{and}\lim_{|V(s)|\to\infty}\overline{\mathbf{h}_{C(s)}}=\boldsymbol{\mu}\quad\text{(a.s.)}$ (1)
where $\boldsymbol{\mu}=\mathbb{E}[\mathbf{h}_{C(s)}]$ is the true mean of the subpopulation neural activity.
To do this, we first compute the mean population response of signal $s$ by averaging over the token-specific hidden state representations $\overline{\mathbf{h}}=\frac{\sum_{j\in V(s)}\mathbf{h}_{(j)}}{|V(s)|}$. A neuron $i$ is hypothesized to be inside C(s){C(s)} if its activity at the time of the concept-specific index fluctuates within a pre-set tolerance level around the mean signal-relevant activity $C(s)=\{i\in W:|h_{i,j}-\overline{h_{i}}|\leq\text{tol}\quad\forall j\in V(s)\}$
@@
and evaluate its quality as a neural population activity classifier for the occurrence of the signal $s$ using true positive rate (TPR) and false positive rate (FPR).
@@ 4.1 Evaluating Reflection Hypothesis on Simple RNNs
First, we assessed the reflection hypothesis using a simple recurrent neural network (Figure 1a and section˜D.1) trained on synthetic sequences containing a known pattern. We begin with sequences in which ABCD appear periodically (Figure 1b), then gradually increase complexity: first by embedding ABCD sparsely within a background of repeated Es (Figure 1c), and then by placing ABCD amid randomly occurring background symbols (E, F, G) to simulate noise (Figure 1d). In each case, the RNN is trained to predict the next character. Across all conditions, the network’s hidden states consistently exhibit activity patterns that reflect the pattern occurrences in the sequence.
@@
We apply Discrete Sequence Chunking to transform continuous hidden unit activities to a symbolic and discretized description. We denote population activities by the alternation of indices that mark the belonging cluster centroid of individual neurons and visualize the population activity by the indices of the assigned nearest cluster. This symbolic description of the neural activities allows decoding the trajectory of the hidden state and its corresponding input via a look-up table (section˜D.2), and reaches a perfect decoding accuracy on the test set. From the symbolic description of neural trajectories, we can then apply chunk discovery methods to learn a dictionary of symbolized chunks. This dictionary contains the maximally recurring patterns inside the network, which consistently reflect the patterns in the input.
@@
We then test whether the extracted chunks causally alter the network’s output in a predictable way. Starting from a state predicting B from input A (with prior input E), we replace the hidden state with those corresponding to A, B, or C, while pairing with inputs B, C, or D, reliably shifts the output predictably to C, D, or E (Figure 1 a).